40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1382     Entry time: Tue Mar 10 04:55:41 2009
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: Locking: 3.7kHz large oscillation 
Yoichi, Jenne, Alberto,

As I reported on the last Thursday, there is a large oscillation in CARM and DARM error signals (attm1).
I put notch filters (3.75kHz, Q=10, 30dB) in the CARM and DARM loops. This let us go up to the arm power of more than 20 and stay there for a while.
The dashed curves in the attm1 are the spectra when the notches are off, and the solid curves are when the notches are used.
We could somewhat suppress the DARM peaks but not CARM.
Of course this is clearly not a good solution. We should find the cause of the oscillation and kill it.

Attm2 is the spectrum of the PO_DC signal flowing in the CM board measured by the SR785. More specifically, CH1 is TP1A and CH2 is TP2A of the CM board.
This was taken right after the AO path was engaged. At this stage, the AO path gain is very low. But you can already see a seed of the oscillation in the spectrum.

Attm3 shows the same spectra taken after the arm power is increased to 4 but before the PO_DC hand off. You can see large peaks around 3.75kHz.
After this, the peaks grow as the power goes up.

Attm4 is the loop gain of the AO path after the PO_DC hand off (arm power = 4).
Attm5 is the zoom of the same TF around 3.7kHz. Clearly there is something wrong at this frequency. We should check the CM board and the MC board as well as the SPOB PD.

One time I was able to go up to arm power = 27 or so. At this power level, the DARM loop started to oscillate, probably, around the UGF.
However because of the 3.7kHz problem, we can't stay at this power level long enough to make diagnostic measurements (like open loop TF).
We should tackle the 3.7kHz issue first.
Attachment 1: CARM_DARM_Spectra.pdf  40 kB  | Hide | Hide all
CARM_DARM_Spectra.pdf
Attachment 2: PODC_Spe_AOPath_Engaged.png  42 kB  | Hide | Hide all
PODC_Spe_AOPath_Engaged.png
Attachment 3: PODC_Spe_before_PODC_handoff.png  43 kB  | Hide | Hide all
PODC_Spe_before_PODC_handoff.png
Attachment 4: AOGain3.png  50 kB  | Hide | Hide all
AOGain3.png
Attachment 5: AOGain2.png  45 kB  | Hide | Hide all
AOGain2.png
ELOG V3.1.3-